OpenStudy (anonymous):

how do i solve for x in √2x-4) - (√x-5) = 3

OpenStudy (jamesj):

\[ \sqrt{2x - 4} - \sqrt{x-5} = 3 \] Yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

HI JAMES!

OpenStudy (jamesj):

Write u = sqrt(x-5). Then x = u^2 + 5 and sqrt(2x - 4) = sqrt(2u^2 + 6) Hence we can rewrite the equation above as sqrt(2u^2 + 6) = u + 3 Now square both sides: 2u^2 + 6 = u^2 + 6u + 9 Hence u^2 - 6u - 3 = 0 Solving for u, u = 3 +- 2sqrt(3) , or u^2 = 21 + 12sqrt(3) ONLY, because the other solution is negative Thus x = u^2 + 5 = 26 + 12sqrt(3)

OpenStudy (anonymous):

the answers are 6, and 30

OpenStudy (jamesj):

No it's not. For example, substitute 6 back into the equation and you get sqrt(12 - 4) - sqrt(1) = sqrt(8) - 1 which does not equal 3.

OpenStudy (jamesj):

However, 6 and 30 are solutions of sqrt(2x + 4) - sqrt(x - 5) = 3. Notice the change from - to + in the first sqrt. !!!!!!!

OpenStudy (jamesj):

You can show it from scratch by following the procedure I just used for your original version of the problem. I think I deserve a 'good answer' for that!

OpenStudy (anonymous):

so can i solve sqrt(2x + 4) - sqrt(x - 5) = 3.

OpenStudy (jamesj):

Yes, make the same substitution I did for u and follow it through, step by step. What grade are you in?

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Hi james !!

OpenStudy (anonymous):

go to chat